North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6

Thursday 31st August 2017

After a day varnishing the tender, and unexpectedly meeting Bill Dobson whose canal trip arrangements had been disrupted, at 16:30 the wheel sets arrived.

Over there on the far line - Nigel Hall

Just about down - Nigel Hall

The intent had been to cover the wheelsets with polythene sheeting after unloading to keep them dry for painting. However, whilst the rest of the wheelsets were being unloaded there was a heavy thunderstorm. There didn't seem much point in covering to keep the water in.


Wednesday 30th August 2017

Ian was a late comer yesterday, arriving about twelve o' clock just in time for lunch, to find me cleaning the cylinder cladding sheets (having spent much of the morning chatting to the public and MPD staff!) and Steve Hyman drilling out the sheared cladding bolts, with Nigel Hall prepping the tender for varnishing.

After lunch Ian removed the small cotters so as to remove the piston rod main cotter. Steve and I carried on with our respective jobs. Ian managed to remove both small cotters but only removed the L/H main cotter in spite of a serious attack on the one on the right. The L/H one is now in the workshop with its small cotter attached. Both cylinder cladding sheets are now cleaned and stored in Deviation shed ready for sanding down and painting .

Only one of the sheared bolts has been drilled out. This took all day as they are extremely hard. Nigel completed varnishing the R/H side of the tender. He intends completing the L/H side today (Thursday).

If any one is available on Saturday there's a few jobs to do - the remaining small cladding bolt to drill out of the R/H cylinder casting, pistons to remove (for which the splitter is required from Darlington, if it's there?), the steam brake to remove and examine, clean and paint the wheelsets, and R/H main cotter to remove just for starters. Ian looks forward to seeing you on Saturday. There's plenty of in date milk in the fridge, but the biscuit stocks could do with replenishing.

Elsewhere, the wheel sets were delayed on their return from South Devon because of a problem with the back load from the North back to the South West. That has been resolved though and the wheel sets are now due back at Grosmont late this afternoon. The MPD has been asked to make sure they are sheeted up so that they are dry for cleaning and painting over the next few working parties. This will also enable the MPD to re-start work on the axleboxes, which they estimate will take about one month to complete.

Mark O'Brien was on with tapping the top set of flues yesterday. He has now found that he can deal with the ones that were fouled by the crown stays by tapping from the smokebox end so he can get the job finished without having to take any drastic action with the crown stays. Because of absences today and tomorrow he is not able to do any more this week, but hopes to have the job finished by the end of next week at the latest, so that the final bottle ends can be ordered. A plan for re-ending the bottle ends is also being finalised.


Saturday 26th August 2017

Bill Dobson becomes Q6 Locomotive Engineering Manager

Today was a NELPG / NYMR meeting held in the MIC room at Grosmont to discus and plan the rebuild of the Q6. This was chaired by our new Q6 caretaker "Bill Dobson," with NYMR shed manager" Paul Middleton" in attendance. Minutes taken by Colin Foxton. The meeting was well attended by NELPG members young & old and lasted nearly two hours. Much was discussed how we are to get loco back into traffic with help from the NYMR and also our own members.

And then to Work

Nigel Bill and myself removed the cylinder cladding sheets to examine the stitching of the cylinder which was done several years ago this was found on the R/H cylinder in two places and all seems ok. In the process we sheered two small cladding bolts which will have to be drilled out we did try but our drills were blunt , these still to do. The next job is to remove the pistons.

Nigel Hall continues with the tender, priming all the chips in the paint work . Andy took Bill around deviation and Q6 pointing out where parts are stored and checking on the progress of work done up to now. Gordon Wells was continuing with the wiring in the pit in Deviation shed.

Andy's wife Joan made us all a welcomed cup of tea also washing our dirty cups up thanks Joan.

And Play

Bryan Orange and the Jv's were having fun on the tandem volocipede on the back road behind Deviation shed.

JVs having fun on the Volocipede - Nigel Hall

More JVs having fun on the Volocipede - Nigel Hall


Wennesday 23rd August 2017

Well we start again after quite a few weeks away from the Q6, although Nigel has been working on the frames priming and glossing.  Today with Chris ,Nigel, Bryan and five JV's. The weather was fine but muggy then about 11.15 it became almost dark we had an almighty thunder storm, lots of visitors sheltering in the north end of the running shed!   We did  achieved a considerable amount of work in preparation to get the locomotive in a forward mode. JV's cleaning inside the frames doing a  great job, two of our JV's were seconded to needle gun duties for the Diesel fitters. Chris cleaning the piston rods which were getting rusty. Both now clean and oiled. Nigel was painting the tender touching up lots of small chips on tender sides with black primer.

I cleaned the flue tube and small tube holes with an emery flap wheel and also the holes in super heater header. This will give Mark more time to get on re- tubing as won't have to be cleaning these holes before he can proceed.  Also a start was mad cleaning the smoke box door seal channel this may need needle gunning on Saturday. Chris and I had lunch in Deviation shed, we are just about out of tea bags but managed by having coffee for elevenses nice change. See you Saturday.


Tuesday 22nd August 2017

Mark O'Brien from the NYMR visited South Devon Engineering last week and took the attached photographs of the Q6 wheelsets which are now completed. They are due to be collected from South Devon next week and delivered to Grosmont on Wednesday 30 August.

The complete driving wheelset at South Devon Railway - Mark O Brien

A wheelset on profiling lathe at South Devon Railway - Mark O Brien

Another complete wheelset at South Devon Railway - Mark O Brien

The dinal complete wheelset at South Devon Railway - Mark O Brien


Wednesday 23rd August 2017

Although there have been no formal Q6 working parties at Grosmont for some weeks now, because of other commitments and work being done by others off site, Nigel Hall has been working quietly away on painting the frames underneath the running plate. That work is now virtually finished, the wheelsets are due back from South Devon at the end of the month, the tapping for the flue tubes in the firebox tubeplate is now just required for the top row, Mark O'Brien was going to order the bottle ends for the 16 tubes that have already been tapped to minimise further delays, and he is pursuing arrangements for welding new ends on the elements. Mark has also identified some preparatory work that needs to be done before the flues are fitted, and a working party, with Ian Pearson as RO, will be held next Wednesday (23rd) at Grosmont to make a start on these tasks.

So if you are looking for a break from the summer holiday, wanting to get away for the day and see the purple heather on the moors, or getting back into the swing of working on the near centenarian Q6, why not come along on Wednesday and give us a hand from 9.30. Tea and biscuits will be available on tap in Deviation Shed, there will be visitors to chat to, and hopefully donations to extract for the collection box, if nothing else. Look forward to seeing you!

Hopefully, working parties will continue on a regular basis from now on, as there will be a lot of work to be done during the next few weeks to get the Q6 back together and hopefully operating before the end of the season. Bill Dobson has now taken on the Q6 caretakership, to ease the workload on Paul Hutchinson, and will be sorting out the future work programme in conjunction with Paul Middleton and the NYMR MPD, so watch this space.


Monday 31st July 2017

Latest photos from the SDR showing one of the intermediate wheels in the profiling lathe. Latest estimate for completion of work is early next week.

Q6 wheel in the profiling lathe - Paul Pridham

Q6 wheel in the profiling lathe - Paul Pridham

Q6 wheel in the profiling lathe - Paul Pridham

Q6 wheel in the profiling lathe - Paul Pridham


Monday 24th July 2017

The tyre and Gibson ring being riveted to a wheel at South Devon Railway - Rob LeChevalier


Friday 14th July 2017

Latest update from the SDR

All tyres fitted. Rivets now in stock for fitting to the tyres. This has been the cause of limited progress this week.

The front wheel set is in the lathe the journals are within 0.007" of true. This is acceptable. The journals and boss faces are being polished there will be a small charge for this.

Harry Frith is going to NDT the wheel sets on the 24th.

By the end of next week we expect good progress to have been made on the riveting of the tyres and profiling of the wheels. We are looking at delivery being in the first week of August.


Wednesday 28th June 2017

A very wet and miserable day at Grosmont today, with Derek, Jon ,Ian Chris and Nigel. Not a great deal to report today, Nigel continuing painting the frames two coats of primer applied to each side. Derek has cleaned all the rubbish from the inside of the drag box and found a spoon and the missing metal ticket from the" Staff & Ticket" section Newbridge - Levisham ,which was lost about two years ago. Jon and I assembled the L/H valve together loosely keys still to make for these valves.

The dome stud hole was re- tapped and a new stud fitted - Ian Pearson

On the dome the stud hole was re- tapped and a new stud fitted see photo. There will be no working parties on Saturday 1/07/17, Wednesday, 5/07/17, Saturday 8/07/17, Wednesday 12/07/17. This is due to holidays and other commitments.


Page 22 of 30